A 57-foot long piece of rope is cut into two pieces so that one piece is three feet shorter than four times the other piece. Find the length of each piece.
Answer by oberobic(2304)

The 57-ft rope is cut into two pieces: x and y.
y is 3 ft shorter than 4x: y = 4x-3
.
x + y = 57
.
substitute
.
x + 4x-3 = 57
.
5x = 60
.
x = 12
.
y = 4x-3
y = 4(12) -3
y = 48-3
y = 45
.
The two pieces are 12 and 45 ft in length.
.
Done.